the length of the fence required to enclose a rectangular field is 3000 meters. What are the dimensions of the field if it is known that the difference between it's length and width is 50 meters
Answer by simranb(63)

Hi,
As you said,
l - b = 50 --------Eqn 1
And the perimeter is the same as the fence required
So,
2(l+b)=3000
l + b = 3000/2
l + b = 1500 ---------Eqn 2
l - b = 50
l = 50 + b
Substitute this value of l into eqn 2
Therefore,
(50 + b) + b = 1500
50 + 2b = 1500
2b = 1500-50
2b = 1450
b = 1450/2
b = 725m
Therefore ,Substitute the value of b into Eqn 1
Then l-b = 50
l-725=50
l = 50+725
l = 775m
Hope you found your answer!]
Simran